OpenStudy (anonymous):

what is the sqrt root of 26^3 and 14c^4

OpenStudy (anonymous):

135.57 \[c^2\sqrt{14}\]

OpenStudy (campbell_st):

\[\sqrt{26^3} = 26^{3/2}\] \[\sqrt{14c^4} = c^2\sqrt{14} = (14c^4)^{1/2}\]

OpenStudy (anonymous):

I dont understand I put both answers in the checkpoint assignment and it says it is incorrect

OpenStudy (anonymous):

132.57 or 2 square root 4394 c^2 square root 14